### v1

# KinesCeTI: A Modular and Size-Adaptable Force Feedback Glove with Interchangeable Actuation for the Index and Thumb

Abstract
Force feedback gloves in haptic applications remain constrained by limited adaptability, simplified feedback and fixed architectures that limit force feedback versatility. To address these challenges, we present KinesCeTI, a modular force feedback exoskeleton for the index and thumb, designed as a multipurpose device adaptable to a wide range of hand sizes. The glove incorporates interchangeable thimbles for fingertip or phalanx attachment and a bidirectional tendon transmission that supports both passive and active feedback. It is combined with a modular actuation design, where different feedback systems may be attached. The system was tested with two actuation modules: a compliant ratchet-pawl braking mechanism for passive feedback and a novel one-way clutch for variable active feedback, newly introduced here. The system was evaluated in three user studies with 20 participants each, assessing ergonomics, actuation performance and usability in both real and virtual tasks. Results indicate that the glove adapts to different hand sizes and provides effective feedback with both mechanisms, highlighting its potential as a versatile platform for haptic research.
footnotetext: This manuscript was submitted to IEEE Transactions on Haptics on October 8, 2025. It has not been peer reviewed. This preprint may differ from the final published version.
I Introduction
Haptic technology has received growing attention in recent decades, following the development of acoustic and visual interfaces throughout the previous century. As their performance reached advanced levels, the integration of touch as an additional modality emerged as a frontier in human-machine interaction. However, despite its progress in recent years [yang2021recent], realistic haptic feedback remains challenging due to technological limits, the complexity of touch and application demands. Haptic systems have been explored in multiple areas, including medicine [licona2019applications], cultural heritage [brewster2017impact], affective touch [eid2015affective], teleoperation [pacchierotti2023cutaneous] and Augmented and Virtual Reality (AR/VR) [tong2023survey]. Its relevance is highlighted in concepts such as the Tactile Internet [fettweis20145g] and the Metaverse [lee2024all].
Because skin sensitivity varies across the body [myles2007tactile], haptic designs are heterogeneous [pacchierotti2017wearable, basdogan2020review, yin2021wearable, culbertson2018haptics, giri2021application]. However, out of all body parts, the hands are the main target for haptic research, due to their central role in manipulation and interaction tasks. Due to their dexterity, sensitivity and force capabilities, numerous types of devices have been developed. These include thimbles, gloves and externally grounded interfaces, each exploring diverse type of stimuli, including pressure, force and vibration [pacchierotti2017wearable]. Nevertheless, no standard design has emerged yet.
This work focuses on a specific category of haptic interfaces: kinesthetic or force feedback gloves. These systems have been extensively researched and developed, with various designs exploring different actuation and sensing technologies, transmission systems and overall structures [tiboni2022soft, du2021review, ferguson2020hand, caeiro2021systematic]. However, as discussed in the following section, current solutions face limitations in feedback capabilities, adaptability to different users and ergonomics.
This paper outlines key design considerations and presents a modular and adaptive hand exoskeleton that implements them. Its actuation potential is demonstrated with two customized force feedback mechanisms developed in this work, which are evaluated alongside the system’s ergonomics and mobility in three user studies. The results are discussed in relation to the glove’s features compared to the state of the art and its potential as a platform for haptics research.
II Force Feedback Gloves: Design Considerations and State of the Art
The design of haptic gloves must begin with the human hand: its size, strength and mobility. Anthropometric studies for healthy subjects report large variability in hand and finger dimensions [chen2011human, mclain2010use, tilley2001measure]. The hand is also capable of considerable force, with mean grip forces up to 55-56 kg and pinch forces around 10 kg [mathiowetz1985grip, angst2010prediction]. Regarding mobility, the human hand has 21 Degrees of Freedom (DOFs) in the fingers and thumb [wang2018toward], which are distributed as follows. Each finger (index, middle, ring, little) includes three joints (metacarpophalangeal (MCP), proximal interphalangeal (PIP) and distal interphalangeal (DIP)) with flexion/extension at every joint and abduction/adduction at the MCP, totaling four DOFs. The thumb consists of three joints (carpometacarpal (CMC), MCP and interphalangeal (IP)) with flexion/extension at each of them and abduction/adduction at the CMC and MCP, totaling five DOFs. Taking into account the hand’s capabilities, the ideal haptic glove would therefore include: high-fidelity force feedback with low latency, accurate motion tracking below perceptual thresholds, ergonomic design, imperceptibility when inactive, adaptability across hand sizes, ease of use, safety and affordability.
From a design perspective, each desired feature imposes new requirements that are often interrelated and sometimes contradictory. To analyze them, we adopt the architecture described by Wang et al. [wang2018toward], which divides exoskeletons into five subsystems: sensing, actuation, transmission, control and structure. Design possibilities for each subsystem are extensive [du2021review, sarac2019design], although interdependent. For instance, the type of control is constrained by sensing and actuation.
Within this framework, we can perform an evaluation of the ideal characteristics, being the most restrictive aspects motion detection and haptics. With regards to feedback modalities, from this point forward this work focuses exclusively on kinesthetic feedback, leaving tactile and thermal modalities outside the scope of this work. Even with this limited approach, motion tracking and force feedback must address the multiple DOFs of the hand. In the context of glove design, DOFs can be categorized in three levels. The first type are passive DOFs, the simplest to implement, as they only imply passive motion. Sensed DOFs add complexity by also measuring movement, requiring electronics and, when needed, markers or attachment elements. And last, feedback DOFs, the most challenging in terms of design, requiring the implementation of actuators, their corresponding force transmission and attachment systems and additional control and electronics. Each type entails increasing demands in terms of weight, volume, power and complexity.
Ideally, every DOF would be both sensed and susceptible to feedback. In practice, the limited space on the hand, the high number of DOFs and technological constraints (particularly in actuation) require trade-offs. Moreover, actuated DOFs are not inherently a binary feature (force vs no-force) but involve various qualitative dimensions, required for achieving complete kinesthetic feedback. Some authors, like Wang et al. [wang2018toward], have analyzed such aspects. In this work, we present an expanded version of different force feedback dimensions:
  • Feedback direction, being some systems capable of feedback in only one direction for a single DOF (e.g., resistance during hand closing), while the ideal system would enable bidirectional feedback. Moreover, bidirectional systems may provide force in both directions simultaneously or at either of them at a particular moment.
  • Transition between states. Being the actuation system capable of switching seamlessly and immediately between the free motion and the feedback states. Such feature requires backdrivability, in order to experience minimal resistance when no force is provided, while enabling immediate feedback when required.
  • Stiffness generation. To recreate different types of objects and interactions, the system should render a wide range of stiffness profiles, from soft objects (e.g., foam) to rigid boundaries (e.g., a glass surface).
  • Type of force. Systems may generate passive force, which resist motion (e.g., friction and brake mechanisms), or active forces, that in addition can induce movement. The latter offers a broader range of feedback capabilities, although at the expense of increasing complexity.
Meeting both these dimensions and the aforementioned overall requirements is highly challenging, since all glove subsystems must support them and many are mutually conflicting. High-fidelity tracking and feedback across all 21 DOFs demand extensive actuation and sensing, yet current technologies make such designs heavy, bulky and power-intensive, reducing wearability and ergonomics. As a general rule, simpler unidirectional passive systems are easier to implement, while bidirectional active designs are far more complex. A further trade-off exists between anatomical adaptability and structural design: soft gloves provide good ergonomics but limited user adjustability, whereas rigid exoskeletons can adapt to different hand sizes at the cost of comfort.
While numerous design strategies have been proposed [pacchierotti2017wearable, du2021review, wang2018toward], all force feedback gloves inevitably involve trade-offs, which are best illustrated by the most representative commercial systems. The leading solutions are HaptX G1 [haptx_glove] and Senseglove Nova 2 [senseglove] with soft-glove architectures. Both provide tactile and force feedback, using a tendon based force transmission system to restrict the grasping motion. This approach yields compact and lightweight designs but limits force feedback to one direction. Their actuation technologies differ, with SenseGlove employing magnetic friction brakes, while HaptX uses a pneumatic brake system, delivering in both cases only passive resistance, complemented by tactile cues for immersion. For hand-size adaptability, both rely on different glove sizes.
An alternative to soft gloves are linkage exoskeletons, which can potentially provide bidirectional feedback and adapt to different users at the cost of higher weight and volume. Dexmo is a prominent example, being originally passive [gu2016dexmo] while in its latest version generates active feedback using customized servomotors. SenseGlove’s earlier DK1 also followed a linkage-based structure, but provided exclusively unidirectional feedback. Numerous other prototypes exist, such as Grabity [choi2017grabity], which simulated weights and opposition grip forces and Wolverine [choi2016wolverine], capable of providing high braking forces with low power consumption. These, however, have not been further developed.
Across these examples, several common limitations of current force feedback gloves emerge. Simplified and often unidirectional feedback is favored to preserve compactness and ergonomics. Hand adaptability is typically addressed through detachable gloves or multiple sizes, but in many devices remains limited. Furthermore, the majority of these solutions adopt closed architectures tailored to a single application or feedback profile, reducing versatility and generalizability. An additional overlooked factor is the acoustic perception of the system. Force feedback presents inherent actuator noise that may influence how users perceive both the feedback and the overall interaction and should therefore be considered in the design process.
These limitations highlight the need for a versatile, modular glove architecture that supports multiple feedback modalities, adapts to different hand sizes and explicitly considers acoustic performance. The following section presents such an exoskeleton, designed with interchangeable actuation and thimble modules and a structural design adaptable to different users.
III KinesCeTI: System Design and Actuation Mechanisms
III-A Design goals
The previous section reviewed design aspects and highlighted the limitations of current technologies and approaches. Building on this analysis, we present the KinesCeTI exoskeleton, developed to address these challenges. It’s designed based on the following main goals:
  • Adaptable to a wide range of adult hand sizes.
  • Unhindered hand motion with high backdrivability, enabling force feedback only when required.
  • General purpose force feedback device for haptic applications, with especial focus on AR/VR. Thus, must be possible to implement different types of force feedback.
  • Accessible in terms of materials and components.
III-B Overall structure
After establishing these design decisions, we proceed to choose the subsystems that enable them. The first consideration is the type of structure and transmission system. Although soft actuation systems have advanced significantly in the last decade, they still lack the maturity and accessibility required for bidirectional systems [tiboni2022soft]. Consequently, a linkage-based architecture placed on the dorsal of the hand is chosen, being both wearable and portable. Because it is grounded to the hand itself, kinesthetic feedback can be applied only to the fingers relative to the hand, but not to the hand as a whole.
One important design decision is the manufacturing method. To ensure low cost and accessibility, most components are made of common polymer-based materials (PLA, PETG, TPU), with the exception of commercial structural parts (e.g., bearings, bolts) and electronics. The exoskeleton is produced through rapid prototyping with additive manufacturing, specifically using Prusa MK3S+ and Prusa Mini+ 3D printers.
In order to focus on developing and testing an effective design, this exoskeleton is designed for the index finger and the thumb, as together they represent all fundamental motion types of the human fingers. Since the index is anatomically similar to the middle, ring and little fingers, its design can be adapted to those in future iterations. In contrast the thumb has a unique anatomy and kinematics, requiring a different mechanical design.
A linkage architecture is implemented for both fingers. The base structure consists of a rigid dorsal platform, attached to the hand with two adjustable strap bands with buckles: one around the knuckles and the other around the wrist. For ergonomic purposes, its contact area with the hand consists of a dorsal-shaped TPU structure with a layer of neoprene rubber for comfortable skin contact. The electronics hub, referred to as the Main Hub, is mounted on this platform. All actuators and sensors are connected to it throughout different connection modules distributed across the exoskeleton. For hand orientation tracking, a 9-DOF IMU (Adafruit BNO055) is embedded within the Main Hub.
The index structure is adjacent to the Main Hub and incorporates a revolute joint that enables abduction/adduction of the index, with its angle measured by a contactless magnetic rotary encoder (AS5600), as shown in Fig. 6. Mounted above are the Knuckle Hub with the linkage and transmission systems, which are nearly identical for both index and thumb, differing only in linkage lengths (explained later). An overview of the exoskeleton is shown in Fig. 1. As the thumb’s ab/adduction produces a wide motion range, the base structure must adapt accordingly. To this end, a structural extension is designed for the thumb oriented at 120 from the index finger in the coronal plane and tilted 30 downward in the sagittal plane (plane conventions follow [bullock2012assessing]).
Figure 1: Overview of the KinesCeTI exoskeleton, showing dorsal placement, Main Hub and attachment system.
To enable unrestricted motion of the thumb, the base structure incorporates two sensed and actuated DOFs. This system, shown in Fig. 2(c), consists of two concentric shafts, with their angles measured by two AS5600 on the outer structure. The first sensed DOF of rotation is for the overall shaft and referred to as simple ab/adduction, allowing up to 90 of motion. The second sensed DOF, combined ab/adduction, includes both this first motion and the ab/adduction of the Knuckle Hub on top of it, transmitted through a bevel gear mechanism. With this bevel mechanism, the thumb’s Knuckle Hub has a range of motion of $\pm 90^{\circ}$ (180 total), with a gear ratio of 13:19 between the Knuckle Hub and the main shaft.
All rotational systems have mechanical stops at the end of their motion range to ensure safe operation. Both shafts also receive feedback from customized ratchet-pawl mechanisms (Fig. 2(a)-(b)), described later in the actuation section. In addition, a compression spring applies a small pulling force to the mobile thumb structure when fully extended, reducing inertia and facilitating user motion.
Figure 2: Thumb structure: (a) frontal view and ratchet-pawl mechanism for the bevel system, (b) rear view and ratchet-pawl mechanism for the simple ab/adduction shaft and (c) side view highlighting the main components.
III-C Linkage system
In both fingers, an assembly consisting of a Knuckle Hub, housing both actuators and sensors and the tendon-linkage system, is mounted on the abduction/adduction shaft. The assemblies are almost identical, differing only in linkage lengths, whose calculation is explained in this subsection. Correct sizing is necessary in order to enable adaptation to different hand sizes. Numerous publications describe the variability of human hand motion and dimensions, using in this work various studies as references [chen2011human, mclain2010use, ueki2010development, garrett1971adult, ozsoyattempt, ash1996proximal, kong2017investigation, jones2006human]. However, it should be noted that these datasets do not represent the entire human population, as they are limited to specific countries, age groups and professions.
The design objective is therefore to create a structure that adapts to most human hands across their full range of motion. To this end, a 2D geometrical analysis is conducted, where both fingers and linkages are modeled as lines without thickness. The chosen structure consists of two linkages in series with revolute joints, with the first joint at the Knuckle Hub and the last at the thimble. Regarding their geometries, the first linkage is a straight beam, while the second is optionally arched. Input parameters include three phalanx sizes (small, medium, large), finger joint angles (tested in 10° steps) and thimble attachment either at the fingertip or the second phalanx, with 20 mm offset perpendicular to their middle points. Values are shown in Tables I and II. Linkage combinations are tested for lengths of 60–200 mm and 10 mm steps for both links. Each combination is evaluated for coverage of the motion range and potential collisions with the finger. When these occur, an arched second linkage (up to 30 mm height) is tested. Combinations failing both checks are rejected. The overall testing process is shown in Fig. 3.
TABLE I: Finger Joint Motion Ranges and Start Joint Coordinates
Finger MCP/CMC PIP/MCP DIP/IP Start joint
Index $90^{\circ} 0$$110^{\circ}$ $90^{\circ} [0$, 45][0,\,45] $m$
Thumb $70^{\circ} 0$$80^{\circ}$ $90^{\circ} [0$, 70][0,\,70] $m$
TABLE II: Phalanx Length Sizes ($m$)
Size Finger Phalanx 1 Phalanx 2 Phalanx 3
Small Index 35.4 14.0 10.8
Thumb 57.0 10.0 22.0
Medium Index 58.36 21.75 26.76
Thumb 78.57 20.02 32.40
Large Index 100.0 32.0 36.0
Thumb 110.0 31.0 45.0
Figure 3: Linkage algorithm: (a) successful configuration with variables indicated, (b) invalid configuration due to insufficient linkage length, (c) invalid configuration caused by intersection with the finger at fingertip attachment and (d) valid configuration using a modified arched geometry for the second linkage.
Figure 4: Algorithm results for (a) index and (b) thumb. Green tiles mark successful combinations, with values indicating the arched height.
Among feasible solutions (Fig. 4), the shortest configurations are selected to minimize weight and volume: L1 = 170 mm, L2 = 130 mm and 30 mm height for the index and L1 = 160 mm, L2 = 90 mm and 20 mm height for the thumb. To preliminarily evaluate force capabilities, load tests with perpendicular forces were applied to the linkages. For the first linkage, only the index configuration (L1 = 170 mm, corresponding to the longest case) was tested, withstanding a maximum load of 52.6 N. For the second linkage, the index configuration (L2 = 130 mm) withstood 45.3 N, while the thumb configuration (L2 = 90 mm) sustained 62.5 N. These tests were performed without additional structural components (bolts, pulleys, or tendons) and represent preliminary feasibility checks rather than definitive strength measurements.
III-D Modular thimbles
The second linkage terminates in the thimble assembly, which serves as the end effector and provides two revolute joints: one for flexion/extension and one for abduction/adduction. Although the latter is already present at the knuckle joint, this extra DOF is required to maintain alignment with the finger’s axis, compensating for the plane misalignment between finger motion and linkage orientation. The thimble assembly includes an attachment system consisting of a slider–lock mechanism, with an insertion slot and a compliant PETG latch that secures it in place. This system is shown in Fig. 5(a) and its operation in Fig. 5(b).
Two types of thimbles were developed. The first is designed for fingertip attachment, which requires customized shapes because the fingertip is a free end lacking a joint to secure the thimble and exhibits greater variability in geometry (Fig. 5(c). For flexibility, the chosen material is TPU 95A. These thimbles are produced in ten sizes for each finger with inner diameters from 14 to 23 mm (1 mm step). Thumb thimbles include additional width variations (2-6.5 mm extra), resulting in a total of 20 designs. All designs are illustrated in Fig. 5(e). The second type is a single-size module for the second phalanx, consisting of a polypropylene strap secured by a PETG hub with an integrated spring latch, which releases when pressed (Fig. 5(d)).
To evaluate the force capabilities of the thimble assembly, preliminary load tests were conducted on both types. For the strap design, five samples were tested, withstanding loads of $98.98\pm 14.55$ N before failure, which consisted of strap loosening while remaining operational. For the TPU fingertip thimbles ten samples were tested, withstanding $45.42\pm 6.11$ N. In this case, failure occurred when the TPU slider pulled out of the PETG holder, after which the thimble remained functional.
Figure 5: Thimble system: (a) compliant latch design and placement, (b) latch operation, (c) TPU fingertip thimble and attachment, (d) strap thimble and attachment and (e) TPU fingertip thimble designs for index and thumb.
III-E Force transmission system
The linkages, thimbles and base structure form a continuous mechanical chain connecting the fingers to the dorsal side of the hand. However, this system is mechanically passive and therefore requires additional components for sensing and force feedback. To minimize weight and inertia on the fingers, both actuation and sensing elements are placed at the Knuckle Hubs. Motion and force are transmitted through a bidirectional closed-loop tendon-pulley architecture, also known as a cable-routed pulley (CPR) configuration with endless cables [grosu2018driving]. In the current design, two tendons made of braided polyethylene line (Spiderwire Dura-4 Braid, 0.35 mm diameter, 35 Kg rated load) are used. Each tendon is anchored at two points: the ratchet-pawl pretensioner on the second linkage, where pretensioning occurs and the main pulley at the Knuckle Hub. The tendon is secured at both ends with constrictor knots, which are reinforced with multiple loops to ensure reliability.
Figure 6: Side view of the index linkage showing the force transmission system and enabled DOFs.
The tendon routing is designed to ensure that the close loop system transmits the angle of both linkages to the main pulley, while intermediate pulleys guide the tendons within the structure and prevent slack, as illustrated in Fig. 6. For sensing, two AS5600 encoders are used: one at the first joint to measure the angle of the first linkage (L1) and another at the main pulley to capture the combined angle of both linkages, thus providing two measured DOFs.
III-F Actuation system
[…]
III-F1 Ratchet pawl mechanism
The first actuation system is a custom compliant ratchet pawl-mechanism (Fig. 7(b)). The ratchet is fixed to the main pulley, while the pawl with a built-in spring is attached to the servomotor. This mechanism provides binary feedback: the pulley is either free to rotate or blocked by the pawl when actuated by the servomotor. It thus enables unidirectional stiff boundary force feedback with high backdrivability. Stiffness is inherently high due to its ratchet-pawl design, although slightly reduced by the compliance of the mechanical elements in series in the actuation chain.
A trade-off exists between resolution, robustness and speed: more teeth increase angular resolution and clutching speed, but reduce robustness. 3D printing constraints also limit possible geometries. In our design, the ratchet had a maximum outer diameter of 16 mm and an inner diameter of 12–13 mm (tooth thickness 1.5-2 mm). Among the tested options (12–30 teeth), the 20-tooth configuration ($18^{\circ}$ resolution/tooth) was selected as the best balance of resolution and strength, with preliminary load tests on three samples yielding relatively high failure loads of 155.04 N, 157.1 N and 131.8 N, consistently above the maximum thimble loads.
Figure 7: Modular actuation system: (a) main pulley socket, (b) ratchet–pawl mechanism, (c) main pulley without feedback, (d) main pulley with ratchet–pawl mechanism and (e) main pulley with one-way clutch.
The mechanism is lightweight (pawl, ratchet and screw weigh 1.74 g) and simple. In addition to the Knuckle Hub design, modified versions were integrated into the thumb structure to provide abduction–adduction feedback (Fig. 2(a)-(b)). Clutching latency was characterized for each subsystem, with 20 samples collected per motor. For the Knuckle Hubs, it yielded a mean latency of $64.475\pm 22.99$ ms (range: 9–108 ms) for the combined 40 samples. For the thumb base, the fixed servomotor showed a latency of $58.5\pm 27.98$ ms (range: 9–108 ms), while the bevel servomotor exhibited $74.25\pm 28.96$ ms (range: 26–121 ms).
III-F2 One-way clutch mechanism
To enable active unidirectional feedback with variable impedance while maintaining good backdrivability, a novel clutch mechanism is developed. In this design, both clutching and feedback are driven by the same servomotor, using an architecture similar to [romeo2021passively] but based on a compliant ratchet–pawl principle. The components, shown in Fig. 8(a), are as follows:
  • Output Ratchet Shaft (ORS): fixed to the main pulley socket, includes a ratchet.
  • Pawl Carrier (PC): surrounds the ORS and houses a compliant pawl.
  • Driving Drum (DD): encloses both the PC and ORS and it’s concentric to them via a bearing and a screw. In this design, it is driven by the servomotor through a direct drive gear. Its inner profile has an irregular geometry that engages the pawl.
  • Friction Ring (FR): provides low friction between the PC and the outer structure (Knuckle Hub).
Figure 8: One-way clutch mechanism: (a) frontal and exploded view of the assembly, (b) clutching operation.
All parts are 3D-printed in PETG, except the FR in TPU. The operation of the clutch, depicted in Fig. 8(b), is as follows:
  • Free motion: ORS and main pulley rotate freely, decoupled from the motor.
  • Clutching: when feedback is required, the motor drives the DD while the FR holds the PC. The inner profile of the DD has a decreasing radius in the pawl area, which pushes the pawl inward to engage the ratchet.
  • Feedback: once engaged, the system enables active feedback, capable of applying resistance or active forces against the finger.
  • Unclutching: reverse motor rotation or opposite finger motion disengages the pawl, restoring free motion.
The clutching angle ($\theta_{\text{clutch}}$) is defined as:
$\theta_{\text{clutch}}=\theta_{\text{pawl}}+\theta_{\text{align}}$ (1)
Here, $\theta_{\text{pawl}}$ is the rotation required for the DD to push the pawl inward, while $\theta_{\text{align}}$ is the additional angle needed to align with the ratchet teeth. $\theta_{\text{align}}$ ranges from $0^{\circ}$ (perfect alignment) to one ratchet step. In our current design, $\theta_{\text{pawl}}$ is approximately $30^{\circ}$ and since the ratchet has 12 teeth, the ratchet step angle is $30^{\circ}$. Thus, the total clutching angle can theoretically range from $30^{\circ}$ to $60^{\circ}$. Clutching latency also depends on finger motion: when the finger is moving, $\theta_{\text{align}}$ decreases, thereby reducing the time required to engage.
The clutch mechanism was implemented in the exoskeleton and measured in both the index and thumb Knuckle Hubs, with 20 samples collected in each. The exoskeleton was kept still during the measurements and was not attached to any hand, to avoid finger movements that may affect the latency. Times were measured between the sent command and motion detection using the corresponding AS5600 sensors. For the combined data, mean latency was $88.7\pm 24.68$ ms (range: 17–152 ms) and mean clutching angles were $52.27\pm 14.57^{\circ}$ (range: 19.46–87.12). The complete clutch module, including bearing and bolt, weighs approximately 4.39 g and measures $\diameter 18.6\times 14$ mm (excluding the servomotor).
III-G Control system
The control system uses an ESP32 in the Main Hub to drive the servomotors and acquire sensor data. AS5600 encoders provide analog PWM outputs that are read through voltage dividers, while servomotor potentiometer signals are buffered with voltage followers to ensure stable readings. Sensing and control electronics are powered via USB from the host computer, whereas the four servomotors are powered separately at 6 VDC (up to 1 A) using a Voltcraft SNG-1000-OC. During experiments, the ESP32 handled sensing and motor control while communicating over serial with a CHAI3D simulation [conti2003chai] running on the host computer, which computes interactions with virtual objects. Further details are presented in the next section.
III-H Overall summary
This section presented the design of a modular and adaptive exoskeleton weighing approximately 376 g, including TPU thimbles and servomotors but excluding the previously described actuation mechanisms. The system integrates interchangeable actuation modules, modular thimbles and an adaptive linkage structure to fit different hand sizes. Its sensing and actuation architecture enables both binary and variable force feedback. These features are evaluated in the user studies described in the next section.
IV User studies
The exoskeleton and its subsystems are evaluated in three user studies:
  • An ergonomics and Range of Motion (RoM) study.
  • A pick and place manipulation task in VR.
  • A pinching softness discrimination task in VR.
Each study was done with written informed consent of each participant. These studies followed ethical guidelines and were approved by the TU Dresden Ethics Committee, with approval number SR-EK-288072024.
IV-A Ergonomics and Range of Motion (RoM) study
IV-A1 Experiment definition
This experiment evaluated the exoskeleton’s adaptability to different hand sizes and mobility range, along with the impact of its structure (without feedback) on real-object manipulation. Two hypotheses were tested: (1) the exoskeleton adapts to a wide range of hand sizes while preserving natural finger motion and (2) the strap thimble allows effective manipulation of real objects.
The evaluation consisted of four parts. First, the measurement of 25 right-hand parameters (Fig. 9, Table III). A subset of these parameters was compared to anthropometric data from the ANSUR database of 1000 males and 1300 females [mclain2010use]. Second, participants performed 16 hand poses with both the free hand and with the exoskeleton using fingertip and strap thimbles (Fig. 10). Third, a customized Six-Hole Peg test, adapted from the Nine-Hole Peg Test [feys2017nine], was conducted to assess fine-motor manipulation (Fig. 11). Participants performed three repetitions of the task with the free hand and with the exoskeleton attached at the second phalanges using strap thimbles, after a short training period. Number of dropped pegs and executing time were measured. Finally, ergonomic evaluation was performed using NASA-TLX questionnaires [hart2006nasa] for both conditions of the peg test and a customized set of 16 ergonomic questions (Table IV).
TABLE III: Hand Measurement Definitions
ID Description ID Description
1 IN-DP length 15 IN-DP circumference
2 IN-MP length 16 IN-DIP circumference
3 IN-PP length 17 IN-MP circumference
4 IN-MC length 18 IN-PIP circumference
5 IN phalanx-crease length 19 IN-PP circumference
6 MI-MC length 20 TH-DP circumference
7 RI-MC length 21 TH-IP circumference
8 LI-MC length 22 TH-PP circumference
9 TH-DP length 23 Knuckles section circumference
10 TH-PP length 24 Palm section circumference
11 TH-MC length 25 Wrist section circumference
12 Knuckles section length 26 IN length to knuckle*
13 Palm section length 27 IN length to crease*
14 Wrist section length 28 TH length to wrist*
Abbreviations: IN=index, TH=thumb, MI=middle, RI=ring, LI=little, DP=distal phalanx, MP=middle phalanx, PP=proximal phalanx, MC=metacarpal, MCP/PIP/DIP/IP=standard joints. *Derived segments, calculated as in Fig. 9.
Figure 9: Dimensions of the hand, according to Table III.
Figure 10: Hand postures performed in the RoM test.
Figure 11: Setup for the Six-Hole Peg task.
TABLE IV: Customized Ergonomics Questionnaire for the first experiment.
ID Question
1 How well does the glove fit your hand?
2 How well does the tip thimble fit on your index finger?
3 How well does the tip thimble fit on your thumb?
4 How well does the strap thimble fit on the second digit of your index?
5 How well does the strap thimble fit on the second digit of your thumb?
6 How would you rate the weight of the exoskeleton?
7 To what extent did the exoskeleton restrict your motion in the index?
8 To what extent did the exoskeleton restrict your motion in the thumb?
9 To what extent did the exoskeleton restrict overall hand motion?
10 How stable did the exoskeleton feel while moving your fingertip thimbles?
11 How stable did the exoskeleton feel while moving your strap thimbles?
12 How comfortable was the exoskeleton material on your skin?
13 Did your hand feel hot or sweaty while wearing the exoskeleton?
14 To what extent did the exoskeleton affect performance in the Peg Test?
15 How would you compare your speed in the Peg Test with and without the exoskeleton?
16 How would you rate the overall comfort of the exoskeleton?
Scales: 0–100 with anchors specific to each question. Intermediate anchors (25, 75) were also shown (e.g., “Somewhat uncomfortable,” “Mostly stable”) Q1–5: 0 = No fit, 50 = Acceptable, 100 = Perfect; Q6: 0 = Very light, 50 = Neutral, 100 = Very heavy; Q7–9: 0 = No restriction, 50 = Some restriction, 100 = Extreme restriction; Q10–11: 0 = Very unstable, 50 = Neutral, 100 = Fully stable; Q12 & Q16: 0 = Very uncomfortable, 50 = Neutral, 100 = Very comfortable; Q13-14: 0 = No/None, 50 = Moderately, 100 = Extremely; Q15: 0 = Much faster free hand, 50 = Same, 100 = Much faster exo.
IV-A2 Experiment results
The study included 20 participants (19 males, 1 female), aged 24 to 44 years (mean= 32.05). Seventeen participants were right-handed and three left-handed. Table V summarizes the hand measurements obtained in our study. To assess generalizability, these results were compared with anthropometric statistics from the ANSUR database [mclain2010use]. Because some of our measurements were not directly compatible due to different measurement definitions, only the overlapping variables were extracted and compared (Table VI). As our dataset includes mixed genders while ANSUR reports separate male and female distributions, we combined the latter into a unified distribution under the assumption of normality. For each parameter, the percentiles corresponding to our minimum and maximum values were computed, yielding an estimate of the coverage (%) of the ANSUR population. Despite the limited set of comparable parameters, the results indicate that our hand measurements span a substantial portion of the ANSUR size range, with an average coverage of $76.01\pm 10.55\%$ across the compared variables.
TABLE V: Measured hand parameters
ID Mean SD Min Max ID Mean SD Min Max
1 25.04 1.81 21.3 28.58 15 50.5 3.51 44 57
2 25.24 2.8 20.98 32.53 16 54.8 2.8 49 60
3 43.66 3.84 36.47 51.61 17 59.05 2.8 54 64
4 73.38 3.69 66.37 79.66 18 66.5 2.64 62 71
5 25.61 2.35 20.29 30.15 19 69.5 3.15 62 74
6 67.11 4.66 59.1 75.38 20 61.45 3.85 55 69
7 63.07 3.35 57.59 69.13 21 67.25 3.52 61 73
8 63.05 5.43 54.73 76.62 22 69.6 3.29 62 75
9 31.36 2.44 26.6 35.45 23 205.1 11.09 174 221
10 34.91 4.29 28.08 42.3 24 218.65 12.3 189 233
11 50.23 4.5 42.81 59.54 25 174.25 8.73 159 188
12 83.89 3.94 74.32 91.15 26 93.93 7.32 80.31 107.61
13 86.52 4.06 76.97 91.87 27 75.88 5.69 64.34 82.79
14 60.25 3.77 52.01 69.72 28 116.5 8.06 101.88 130.07
Units: mm.
TABLE VI: Common Hand Measurements: Comparison with McLain [mclain2010use]
Combined data from McLain [mclain2010use](mm) Comparison results (%)
ID Mean SD Min Max pMin pMax %
1 26.76 2.62 19 36 1.85 75.64 73.78
2 21.75 2.41 14 32 37.42 100 62.58
9 32.41 3.09 22 45 3.01 83.74 80.72
12 84.24 6.71 66 106 6.97 84.85 77.88
14 60.83 5.86 46 82 6.63 93.54 86.91
16 53.67 3.72 45 65 10.49 95.55 85.06
18 64.33 4.05 56 74 28.26 95.01 66.76
21 67.04 3.53 56 81 12.86 79.73 73.94
23 198.41 16.37 158 247 6.8 91.64 84.81
25 161.36 13.6 130 204 43.12 97.49 54.37
27 72.08 5.51 56 92 8.02 97.4 89.39
% == pMax - pMin.
Figure 12: Motion restrictions for the exoskeleton with each thimble attachment in comparison with the free hand.
In the motion tests with fingertip and strap thimbles, participants were able to perform 12 and 10 out of 16 hand postures, respectively, without restriction (Fig. 12). For the remaining postures, restrictions were observed, differing between the two thimble types. Most limitations occurred near the end of the motion range. With the exception of one participant in each setup who reported moderate restriction for the 11th posture, all others reported only minor restriction. Overall, results indicate that the exoskeleton allows most natural hand postures, though improvements remain possible.
In the Six-Hole Peg test, no failures occurred with the free hand and only four failures in total with the exoskeleton (each by a different participant in the third round). As these were rare events, no statistical analysis was conducted on failure rates. In contrast, execution times differed significantly between conditions. For each subject and condition, the average of the three repetitions was calculated. Since data were not normally distributed, results were analyzed using medians and IQRs across participants. Median task time was $19.99$ s (IQR =15.69=15.6922.7822.78) with the exoskeleton, compared to $14.60$ s (IQR =12.79=12.7916.2116.21) with the free hand. The median paired difference was $4.70$ s, 95% CI [$3.52$, $6.80$]. A one-sided Wilcoxon signed-rank test confirmed longer times with the exoskeleton, $W=210$, $z=3.91$, $p<.001$, with a large effect size ($r=.88$). For descriptive context, mean times were $14.89\pm 2.64$ s for the free hand and $20.48\pm 4.94$ s for the exoskeleton, corresponding to a mean increase of $5.58\pm 2.71$ s, or $36.5\pm 13.5$%. These findings show that use of the exoskeleton without feedback increased execution times and reduced performance efficiency. Nonetheless, aside from the four isolated failures, participants were still able to complete the manipulation task with the exoskeleton.
NASA-TLX scores were compared between free hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the exoskeleton (Fig. 13 and Table VII). After Bonferroni-Holm correction, the exoskeleton condition was rated significantly higher in mental demand, physical demand, effort, frustration and overall workload (all $p<.05$, one-sided). Surprisingly, temporal demand and performance did not show statistically significant differences (p = .89 and .5, respectively), even though objectively measured metrics indicate otherwise.
Figure 13: NASA-TLX results for the Six-Hole Peg task.
TABLE VII: NASA-TLX scores of the Six-Hole Peg test reported as median [IQR] and Wilcoxon signed-rank results (n=20).
Subscale Free Exo Wilcoxon (W, z, p, r)
Mental Demand 20 [10.0-53.75] 32.5 [20.0-50.0] 113.0, 2.39, .0410, .59
Physical Demand 22.5 [10.0-40.0] 40.0 [28.75-70.0] 166.0, 3.57, .0016, .84
Temporal Demand 57.5 [41.25-80.0] 42.5 [33.75-77.5] 33.0, -1.23, .8966, .33
Performance 20.0 [5.0-53.75] 30.0 [8.75-40.0] 63.5, 0.67, .5081, .18
Effort 25.0 [10.0-50.0] 37.5 [23.75-66.25] 135.5, 2.20, .0458, .51
Frustration 7.5 [5.0-20.0] 20.0 [5.0-36.25] 84.0, 2.71, .0193, .75
NASA-TLX Total 33.33 [25.625-38.75] 36.66 [27.5-49.16] 168.0, 2.93, .0105, .67
TABLE VIII: Customized Ergonomics Questionnaire Results (Median [IQR])
ID Mdn [IQR] ID Mdn [IQR]
1 80 [73–91.5] 9 23.5 [9–25.25]
2 78 [67.5–90] 10 75 [48.75–89.25]
3 76.5 [63–89.25] 11 82.5 [75–90.25]
4 85 [76.75–96] 12 67 [50–81]
5 80 [65.75–93.25] 13 2 [0–28.25]
6 61 [48.25–67] 14 38 [14–64.25]
7 10.5 [5.25–25.25] 15 26 [20–35.25]
8 23 [13.75–41.75] 16 70 [50–84.5]
Values are reported as median [IQR] on 0–100 scales, with anchors defined in Table IV.
The ergonomics questionnaire results (Table VIII) can be grouped into several categories. Fit (Q1–5) was rated positively, with medians between 76.5 and 85, indicating generally good adaptation of the glove and thimbles. Weight (Q6) was judged moderately heavy (Mdn = 61, IQR = 48.25–67), with most values between neutral and heavy. Restriction (Q7–9) was low overall: the index finger showed the least limitation (Mdn = 10.5), followed by the thumb (Mdn = 23), while overall hand restriction remained below 50 for almost all participants. Stability (Q10–11) was consistently high, with both fingertip and strap thimbles rated mostly stable (Mdn $\geq$ 75). Comfort (Q12, Q16) was also positive (Mdn = 67–70), although not maximal. Heat and sweating (Q13) were reported as minimal (Mdn = 2, IQR = 0–28.25). Finally, participants acknowledged an impact on perceived performance in the Six-Hole Peg test, rating completion as faster with the free hand (Q14–15, Mdn = 26–38), consistent with the observed execution times.
Overall, this first test evaluated ergonomics, mobility and the effects of wearing the exoskeleton. The device enabled most movements with minimal restriction in a group of subjects that moderately resembled the general population, while offering satisfactory yet improvable ergonomics. With regard to movement and task performance, the Six-Hole Peg task showed increased execution time when using the exoskeleton, but very low impact on effectiveness (low failure rate).
IV-B Pick and place manipulation task
IV-B1 Experiment definition
The second experiment assessed the influence of force feedback from the exoskeleton’s ratchet–pawl mechanism compared to free-hand control in a virtual pick-and-place task. Participants were required to grasp a virtual ball and place it into one of two target rings (Fig. 14(a)). The task was implemented in CHAI3D. The independent variables were:
  • Hand condition: free hand tracked with the Leap Motion controller, versus exoskeleton, where the Leap Motion provided global hand tracking while the exoskeleton supplied finger sensing and force feedback.
  • Tolerance threshold for failure with two values (0.025 and 0.03). During object pickup, fingertip distance was recorded and the threshold defined the allowable variation. Smaller thresholds increased the likelihood of dropping or over-squeezing the object (thus higher difficulty). These values were set after pilot testing and are not meant to represent real-world dimensions.
  • Target ring: far (Goal 1) versus near (Goal 2) from the subject, both aligned along the Y-axis (forward from the user) and equidistant from the ball’s initial position on the right.
  • Repetitions: Three per stimulus, resulting in 12 samples per hand condition and 24 in total.
The experimental hypothesis was that force feedback from the exoskeleton would reduce execution time and failure rates and yield more favorable subjective ratings, compared to free-hand interaction in virtual manipulation. Data were collected at two levels. Objective metrics included execution time per trial and the number of failures. Subjective metrics included a NASA-TLX questionnaire for each hand condition, followed by a questionnaire on force and acoustics perception and overall preference (Table IX).
Figure 14: Virtual setups for (a) the pick an place test and (b) softness discrimination experiment.
TABLE IX: Questionnaire for the Virtual Pick-and-Place Task
ID Question
1 To what extent was the force feedback helpful during the task?
2 How noticeable were the sounds coming from the exoskeleton actuation system?
3 How distracting were the motor sounds?
4 How would you rate the annoyance of the motor sounds?
5 To what extent were the sounds from the exoskeleton motors helpful during the task?
6 Which condition do you prefer for this type of task: with the free hand or with the exoskeleton?
Scales: 0–100 with anchors specific to each question. Example anchors:
Q1: 0 = Not, 50 = Moderate, 100 = Extremely helpful.
Q2: 0 = Not, 50 = Somewhat, 100 = Very noticeable.
Q3: 0 = None, 50 = Moderate, 100 = Extreme distraction.
Q4: 0 = None, 50 = Moderate, 100 = Extreme annoyance.
Q5: 0 = Not, 50 = Moderate, 100 = Extremely helpful.
Q6: 0 = Strong free hand, 50 = No preference, 100 = Strong exo
Intermediate anchors (25, 75) were also shown to participants.
IV-B2 Experiment results
The study included 20 participants (16 males, 4 females), aged 26–62 years (mean = 32.25), of whom 18 were right-handed and 2 left-handed. For completion times, mean values were computed per participant for each combination of tolerance and goal. As data were not normally distributed, Wilcoxon signed-rank tests were applied under the alternative hypothesis that the free-hand condition would yield longer times. Bonferroni–Holm correction was applied across the four comparisons. As summarized in Table X, all differences remained statistically significant after correction (p=.0002p=.0002–.0068). The free-hand condition consistently resulted in longer completion times ($\tilde{\Delta}\approx$ +0.4–0.6 s), with large effect sizes in all cases.
Across all trials, more failures (defined as dropping or over-pinching the virtual balls) occurred with the free hand (320) than with the exoskeleton (29). As failures per configuration (goal and threshold) were sparse and commonly zero for the exoskeleton, they were aggregated per participant across all trials and compared. The free-hand condition produced substantially more errors (Mdn = 9.5, IQR = 2.0–31.5) than the exoskeleton (Mdn = 1.0, IQR = 0.0–2.25). A one-sided Wilcoxon signed-rank test confirmed this difference, $W=145$, $z=3.29$, $p<.001$, with a large effect size ($r=.79$).
When comparing data between the two tolerance thresholds (0.025 vs. 0.030), no consistent differences were found. Occasional effects were observed in failures, but these were based on sparse and zero-inflated data and are not considered conclusive. Similarly, no statistically significant differences were observed between Goal 1 and Goal 2 after applying Wilcoxon tests with Bonferroni–Holm correction.
TABLE X: Wilcoxon results for completion times (Leap Motion vs Exoskeleton)
Cfg Exo [Mdn, IQR] Leap [Mdn, IQR] $\tilde{\Delta}$ [CI] W z p r
G1–.03 1188 [1018–1577] 1742 [1157–2344] +540 [–6; 959] 179 2.75 .0042 .62
G1–.025 1133 [1079–1400] 1821 [1377–2635] +611 [262; 977] 199 3.50 .0002 .78
G2–.03 1144 [1058–1534] 1885 [1153–2479] +422 [–98;1174] 170 2.42 .0068 .54
G2–.025 1158 [1041–1446] 1982 [1162–2600] +515 [66;1218] 182 2.87 .0040 .64
Values are medians [IQR] in ms. $\tilde{\Delta}$ = median paired difference (Exo–Leap). CI = 95%. Cfg = configuration: G1/G2 = target goal, .025/.03 = tolerance threshold. $p$ values are Bonferroni–Holm corrected.
NASA-TLX scores were compared between free-hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the free hand. Results are summarized in Fig. 15 and Table XI. After Bonferroni-Holm correction, significant differences were observed in Mental Demand ($p=.008$), Performance ($p=.02$), Effort ($p=.01$), Frustration ($p=.02$) and the overall NASA-TLX Total score ($p=.01$), all favoring the exoskeleton. Physical Demand and Temporal Demand did not reach significance ($p=.155$ each). These results suggest that the exoskeleton condition was perceived as less demanding overall, with lower mental effort, frustration and effort ratings, as well as better perceived performance.
Figure 15: NASA-TLX results for the virtual pick and place task.
TABLE XI: NASA-TLX scores of virtual pick and place test reported as median [IQR] and Wilcoxon signed-rank results (n=20).
Subscale Free Exo Wilcoxon (W, z, p, r)
Mental Demand 37.5 [18.75–71.25] 20.0 [13.75–30.0] 114.0, 3.08, .008, .80
Physical Demand 25.0 [13.75–31.25] 20.0 [13.75–35.0] 114.5, 1.27, .155, .30
Temporal Demand 42.5 [25.0–61.25] 35.0 [18.75–46.25] 118.5, 1.46, .155, .34
Performance 47.5 [20.0–71.25] 15.0 [5.0–26.25] 172.5, 2.53, .020, .57
Effort 52.5 [20.0–70.0] 25.0 [13.75–41.25] 153.5, 2.98, .010, .70
Frustration 52.5 [20.0–73.75] 20.0 [10.0–35.0] 159.5, 2.60, .020, .60
NASA-TLX Total 41.25 [32.08–64.58] 25.83 [16.88–35.63] 153.0, 2.93, .010, .69
With regards to the custom questionnaire, participants rated force feedback as helpful (Q1: Mdn = 64.5, IQR = 40.0–75.0)). Sounds were clearly noticeable (Q2: Mdn = 75.0, IQR = 25.8–76.3), but rated low in distraction (Q3: Mdn = 5.5, IQR = 1.5–23.5) and annoyance (Q4: Mdn = 5.5, IQR = 2.8–12.8). The perceived helpfulness of sounds varied considerably across participants (Q5: Mdn = 27.5, IQR = 7.8–70.3). Finally, overall preference strongly favored the exoskeleton (Q6: Mdn = 85.0, IQR = 60.8–97.8), well above the neutral midpoint of 50.
These results suggest that the exoskeleton with force feedback enhanced task performance relative to the free-hand condition, with shorter completion times and fewer failures. Subjective measures were consistent with this outcome: workload ratings (NASA-TLX) were generally lower with the exoskeleton and participants considered force feedback helpful, while motor sounds were noticeable but not distracting or annoying. Notably, overall preference favored the exoskeleton. Together, these results suggest that this system enhances performance in virtual pick-and-place tasks and is well accepted by users.
IV-C Softness discrimination study
IV-C1 Experiment definition
This experiment evaluated the one-way clutch mechanism in a pinching softness discrimination task in VR. Force feedback was rendered by a pseudo-impedance controller that engages the clutch to provide unidirectional active forces on the fingers. The control algorithm running on the ESP32 computed the motor command as
$u_{i}\;=\;\theta_{\mathrm{eq}}\;+\;0.01\,k\,s_{i},$ (2)
where $u_{i}$ is the motor command (servo position), $s_{i}$ is the virtual fingertip penetration into the object, $k$ is the softness parameter (pilot-calibrated), $0.01$ is a fixed gain and $\theta_{\mathrm{eq}}$ is the equilibrium motor position when the clutch engages the main pulley after feedback is activated.
Three softness levels were tested, selected via preliminary calibration: rigid (A), medium-rigid (B) and soft (C).The experimental hypothesis was that the one-way clutch system would render perceptually distinct softness levels, allowing participants to discriminate between them with accuracy above chance. The task consisted of presenting participants with two virtual balls per trial, each with a different softness, as shown in Fig. 14(b). After probing both, participants indicated which one felt softer. All pairwise comparisons (AB, AC, BC) were tested, each repeated three times, yielding nine samples per participant. Ball positions (left/right) and stimulus order were randomized. Prior to the task, participants performed a brief training (2 min) to become familiar with the procedure.
Data collection included both objective and subjective measures. Objective data consisted of discrimination accuracy (number of correct and incorrect softer-ball selections). Subjective data included a short questionnaire on task perception, force feedback and acoustic awareness (Table XII).
TABLE XII: Customized Questionnaire for the Pinching Softness Discrimination Task
ID Question
1 To what extent could you perceive differences in stiffness between the two balls in each trial?
2 How confident were you in identifying the softer ball in each comparison?
3 How often did the two balls feel indistinguishable in terms of stiffness?
4 How noticeable were the sounds coming from the exoskeleton actuation system?
5 To what extent did the motor sounds distract you from the task?
6 How would you rate the annoyance of the motor sounds?
7 To what extent were the sounds from the exoskeleton motors helpful during the task?
Scales: 0–100, anchors specific to each question.
Q1: 0 = None, 50 = Moderate, 100 = Extremely clear.
Q2: 0 = Not confident, 50 = Moderate, 100 = Completely.
Q3: 0 = Never, 50 = Sometimes, 100 = Always same.
Q4: 0 = Not noticeable, 50 = Somewhat, 100 = Very.
Q5: 0 = None, 50 = Moderate, 100 = Extremely distracting.
Q6: 0 = None, 50 = Moderate, 100 = Extremely annoying.
Q7: 0 = Not helpful, 50 = Moderate, 100 = Extremely.
Intermediate anchors (25, 75) were also provided.
IV-C2 Experiment results
The study included 20 participants (18 males, 2 females), aged 24–39 years (mean = 30.15), of whom 14 were right-handed and 6 left-handed. Performance was assessed as correct or incorrect identification of the softer object. Across all trials, the overall discrimination accuracy was high (Mdn = 88.9%, IQR = 77.8–100.0). As the data were not normally distributed, nonparametric analyses were applied. Because responses were binary (success vs. failure), one-sided binomial tests were conducted against the null hypothesis of chance performance (50%), with Bonferroni-Holm correction across the three pairwise comparisons (AB, AC, BC). Results are summarized in Table XIII. Accuracy was statistically significant above chance for all comparisons (all $p<.001$), with large to very large effect sizes (Cohen’s h=0.69h=0.690.990.99), indicating that participants reliably discriminated between the tested softness levels.
TABLE XIII: Binomial test results for softness comparisons
Comparison Correct (k/n) Proportion 95% CI $p$ (Holm) Cohen’s $h$
AB 49/60 0.82 [0.71, 1.00] <.001<.001 0.69 (large)
AC 55/60 0.92 [0.83, 1.00] <.001<.001 0.99 (very large)
BC 50/60 0.83 [0.73, 1.00] <.001<.001 0.73 (large)
In the custom questionnaire, participants generally reported that stiffness differences between the two balls were clearly perceivable (Q1: Mdn = 67, IQR = 48.3–75.0) and expressed moderate to high confidence in identifying the softer ball (Q2: Mdn = 70, IQR = 47.5–85.0). Still, indistinguishability was occasionally reported (Q3: Mdn = 33.5, IQR = 24.8–54.8), indicating that some comparisons remained challenging. Motor sounds were rated as highly noticeable (Q4: Mdn = 75.0, IQR = 58.3–90.3), but distraction (Q5: Mdn = 10.0, IQR = 0.0–18.5) and annoyance (Q6: Mdn = 14.0, IQR = 0.0–23.5) were minimal. Finally, the perceived helpfulness of sounds varied substantially (Q7: Mdn = 25.0, IQR = 9.3–52.0), suggesting that while audible cues were sometimes useful, their contribution was not consistent across participants.
Overall, the results of this study confirm that the exoskeleton with its one-way clutch mechanism rendered perceptually distinct softness levels, that participants discriminated reliably with relatively high accuracy. Subjective ratings indicated moderately high perception of stiffness differences and confidence in judgments, while acoustic feedback was noticeable but not distracting or annoying. Taken together, these findings support the hypothesis that the system provides effective unidirectional force feedback for softness rendering.
V Conclusions
This work presents KinesCeTI, a modular force feedback glove with flexibility in both its thimble attachment system and its actuation capabilities. The glove integrates a bidirectional tendon routing system and a structure adaptable to different hands, enabling the implementation of external actuation modules. Two mechanisms were developed: a ratchet–pawl brake and a novel one-way clutch, first introduced in this work. The system was evaluated in three user studies, confirming ergonomics across hand sizes, functional operation of both mechanisms, the ability to manipulate real and virtual objects and the feasibility of rendering variable impedance.
Compared to existing systems, KinesCeTI offers advantages in various aspects. First, in hand adaptability, through an adjustable linkage structure and strap thimble design, though fingertip thimbles still required different sizes, which could be simplified with an improved strap-based design. Second, in versatility, as the architecture supports multiple force feedback modalities (passive, active, variable, bidirectional), unlike existing solutions such as HaptX and SenseGlove. The type of stimulus is therefore dependent on the externally attached actuator module. In this regard, latency remains an area for improvement, with mean values of 58.5 ms for the ratchet–pawl and 88.7 ms for the clutch, both showing high variability. While the ratchet–pawl mean latency is comparable to or slightly lower than the 62 ms reported for HaptX, the clutch remains higher, particularly when compared to Wolverine (21 ms). Nevertheless, the presented mechanisms are lightweight, compact, affordable and simple. Another distinguishing aspect of this work is the explicit consideration of acoustic perception, addressed both in motor selection and in user studies, a factor rarely considered in prior designs.
The technical tests and user studies also highlighted areas requiring further development: improving force resolution and control, reducing latency and its variability, extending modular and bidirectional actuation to the thumb structure and further optimization of weight, volume and ergonomics. More precise digit-level tracking is also needed to complement the current overall position sensing.
Overall, KinesCeTI provides a versatile platform for haptics research. Its modular design allows different actuators and modalities to be tested within the same platform, avoiding redesigns required by closed architectures. Knuckle Hubs can host alternative actuation modules, while the thimble system allows easy integration of tactile, thermal or multimodal modules. This enables comparative testing of different actuation principles and applications within a single glove.
Acknowledgment
Funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden.
[Uncaptioned image] Pablo Alvarez Romeo received the M.Sc. degree in Electronic Engineering from the University of Zaragoza, Zaragoza, Spain. He is currently a Research Associate with the Chair of Acoustics and Haptics, TU Dresden. His research interests include haptic interfaces and perception, robotics, mechatronics design and development, wearables, additive manufacturing and rapid prototyping.
[Uncaptioned image] Mehmet Ercan Altinsoy received the graduation degree in mechanical engineering from the Technical University of Istanbul, Istanbul, Turkey and the Ph.D. degree in electrical engineering from Ruhr-University Bochum, Bochum, Germany, in 2005. After receiving his Ph.D. degree, he was with HEAD Acoustics as a Consulting Engineer. Since 2006, he has been with TU Dresden, Dresden, Germany, where he is currently a Professor in acoustic and haptic engineering. His research interests include perception-based engineering, vibroacoustics, vehicle acoustics, electroacoustics, haptic interfaces, haptic perception, whole-body vibrations, product sound and vibration design. Dr. Altinsoy was a member of the International Graduate School for Neuroscience with Ruhr-University Bochum. In 2018, he was awarded a Visiting Professorship from Tohoku University, Japan. He is a Lothar-Cremer Medalist of the Acoustical Society of Germany, DEGA. He is the Chairman of the Vehicle NVH Expert Committee of DEGA and one of the core team members of the Cluster of Excellence Centre for Tactile Internet with Human-in-the-Loop.